OpenStudy (anonymous):

plz help me find the solution to each system of equation by the substitution method. 4/7x +2/7y =2 3x+y =13

OpenStudy (ryan123345):

Hey there! I see that you are new to openstudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! I wish you good luck on openstudy and dont forget to read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Hope you have fun!

OpenStudy (ryan123345):

welcome :)

OpenStudy (ryan123345):

Are you new?

OpenStudy (ryan123345):

cool :)

OpenStudy (anonymous):

3x + y = 13 ---> y = 13 - 3x now sub 13 - 3x in for y in the 1st equation...First lets get rid of the fractions 4/7x + 2/7y = 2 (multiply the equation by 7 to get rid of the fraction... (7)4/7x + (7)2/7y = (7)2 4x + 2y = 14 (now sub 13 - 3x for y) 4x + 2(13 - 3x) = 14 (distribute the 2 through the parenthesis) 4x + 26 - 6x = 14 (combine like terms) -2x + 26 = 14 -2x = 14 - 26 -2x = - 12 x = 6 now sub 6 in for x in the 2nd equation 3x + y = 13 3(6) + y = 13 18 + y = 13 y = 13 - 18 y = - 5 check... 4x + 2y = 14 4(6) + 2(-5) = 14 24 - 10 = 14 14 = 14 (correct) ANSWER: x = 6 and y = -5 or (6,-5)

OpenStudy (anonymous):

Do you have any questions ?

OpenStudy (anonymous):

thanks a lot I have more question 4/7x +2/7y=2 3x+y=13

OpenStudy (anonymous):

solve by graphing y=3/4x-2 y=1/3x +2/3

OpenStudy (anonymous):

y-2x-6=0 1/2y-3=x

OpenStudy (anonymous):

please help me as soon as possible